# SGS-3D: High-Fidelity 3D Instance Segmentation via Reliable Semantic Mask Splitting and Growing

### v1

@@ Abstract
Accurate 3D instance segmentation is crucial for high-quality scene understanding in the 3D vision domain. However, 3D instance segmentation based on 2D-to-3D lifting approaches struggle to produce precise instance-level segmentation, due to accumulated errors introduced during the lifting process from ambiguous semantic guidance and insufficient depth constraints. To tackle these challenges, we propose Splitting and Growing reliable Semantic mask for high-fidelity 3D instance segmentation (SGS-3D), a novel ”split-then-grow” framework that first purifies and splits ambiguous lifted masks using geometric primitives, and then grows them into complete instances within the scene. Unlike existing approaches that directly rely on raw lifted masks and sacrifice segmentation accuracy, SGS-3D serves as a training-free refinement method that jointly fuses semantic and geometric information, enabling effective cooperation between the two levels of representation. Specifically, for semantic guidance, we introduce a mask filtering strategy that leverages the co-occurrence of 3D geometry primitives to identify and remove ambiguous masks, thereby ensuring more reliable semantic consistency with the 3D object instances. For the geometric refinement, we construct fine-grained object instances by exploiting both spatial continuity and high-level features, particularly in the case of semantic ambiguity between distinct objects. Experimental results on ScanNet200, ScanNet++, and KITTI-360 demonstrate that SGS-3D substantially improves segmentation accuracy and robustness against inaccurate masks from pre-trained models, yielding high-fidelity object instances while maintaining strong generalization across diverse indoor and outdoor environments. Code is available in the supplementary materials.
@@ Introduction
Arbitrary instances segmentation within 3D scenes constitutes a crucial task in various domains, including autonomous driving, virtual reality, and multi-modality scene understanding. While recent methods (Chen et al. 2021; Vu et al. 2022; Liang et al. 2021; Engelmann et al. 2020; Lu et al. 2023a; Schult et al. 2023) yield impressive segmentation results when trained on specifically annotated datasets (Dai et al. 2017; Armeni et al. 2016; Robert, Raguet, and Landrieu 2023; Yeshwanth et al. 2023; Liao, Xie, and Geiger 2022), their generalization to open-world environments remains limited. In contrast to the inherent difficulties in obtaining extensive 3D annotations, foundation models trained on massive 2D image datasets have shown remarkable performance and strong zero-shot generalization abilities. This observation motivates a promising direction: utilizing 2D pre-trained foundation models for 3D scene perception and interaction.
Currently, two predominant paradigms exist for executing the 3D instance segmentation: feature-based methods and mask-based methods. Feature-based methods (Peng et al. 2023; Takmaz et al. 2023; Xu et al. 2024; Lee, Zhao, and Lee 2024) primarily focus on learning robust feature representations from 3D point clouds or by lifting features from 2D modalities. Instance segmentation is subsequently achieved by classifying, grouping, or decoding these learned features into instance-level masks. However, these methods commonly suffer from inefficient training and error propagation, which arising from ambiguous feature representations stemming from 2D-to-3D lifting or the inherent high-dimensional feature space.
To mitigate these concerns, mask-based methods leverage the semantic masks extracted from input registered images using Segment Anything Model (SAM) (Kirillov et al. 2023) to derive instance-level assignments for each 3D geometric primitive in the 3D scene, consolidating these 3D primitives iteratively based on their projected masks (Yang et al. 2023; Yin et al. 2024; Nguyen et al. 2024). Despite this advancement, the robust propagation of semantic information from these projected masks is often overlooked, with a primary focus on optimizing strategies for merging 3D geometric primitives while neglecting fine-grained geometric cues. This oversight leads to continuous error accumulation during scene segmentation, ultimately yielding imprecise results, particularly for adjacent but separate instances, as il-lustrated in Figure 1. Although some existing methods incorporate depth information from sensors to enhance back-projection accuracy (Guo et al. 2024; Yin et al. 2024), this reliance is problematic for two key reasons: depth sensors inherently struggle on textureless and highly reflective surfaces, and moreover, they are often entirely unavailable in wild scenes, severely compromising segmentation performance.
@@
We first over-segment the point cloud $P$ into a set of $U$ superpoints, $\mathcal{S}=\{s_{u}\}_{u=1}^{U}$, which serve as our fundamental processing primitives. These spatially coherent and geometrically compact groups of points improve computational efficiency and help maintain structural consistency. To adapt to diverse 3D scenarios, we employ method in (Dai et al. 2017) for indoor mesh inputs and (Robert, Raguet, and Landrieu 2023) for outdoor unstructured point clouds. Furthermore, for each superpoint $s_{u}$, we assume access to a pre-computed feature vector $\mathbf{f}_{u}\in\mathbb{R}^{D}$, derived from a pre-trained segmentation model (Schult et al. 2023), which captures its high-dimensional feature.
@@
In contrast to existing methods that rely on ground-truth depth maps from dedicated sensors, we propose an efficient mapping strategy to compute visibility directly from the point cloud and camera parameters using Z-buffering (Ravi et al. 2020). For a given image $I_{t}$ with its associated camera parameters $(\mathbf{K}_{t},\mathbf{T}_{t})$, the entire point cloud $P$ is projected into the camera’s image plane. The transformation $\pi$ maps a homogeneous world point $p_{i}$ to pixel coordinates $(u_{i},v_{i})$ with depth $z_{i}$, according to $z_{i}[u_{i},v_{i},1]^{T}=\mathbf{K}_{t}\mathbf{T}_{t}^{-1}\mathbf{p}_{i}$. To resolve occlusions, a depth buffer $\mathbf{D}_{t}\in\mathbb{R}^{H\times W}$ is initialized with infinite values. The final value at each pixel $(u,v)$ is populated via a highly parallel rasterization process on the GPU, which efficiently solves for the minimum depth over the set of all points projecting to that location:
@@
Mask-based 3D instance segmentation involves generating masks for the target objects based on input images. For each image $I_{t}$ in a scene sequence $I=\{I_{t}\}_{t=1}^{T}$, the objective is to produce a corresponding set of binary masks, $\mathcal{M}_{2D,t}=\{M_{t,k}\}_{k=1}^{K_{t}}$. With the emergence of foundational models like SAM (Kirillov et al. 2023), mask prediction has become highly precise. However, SAM lacks targeted guidance and often segments objects that are not of interest. Therefore, we apply Grounding-DINO (Liu et al. 2024) to extract box prompts from the images with text prompts. The high-confidence box prompts among them are then used to generate the initial mask predictions.
@@
In fact, the binary masks $\mathcal{M}_{2D,t}$ predicted by the trained 2D vision foundation model exhibit discreteness, which often leads to inaccuracies and inconsistencies in object boundary area predictions across different views. This limitation may increase the instability of the semantic guidance of 2D masks. Unlike Maskclustering (Yan et al. 2024), which purely relies on image pixel consistency between views, we leverage viewing visibility of superpoints to exclusively enhance robustness against inaccurate masks.
@@
Existing methods often compromise segmentation quality by either directly unprojecting 2D masks onto the point cloud (Yang et al. 2023; Yan et al. 2024) or by solely using them as a semantic prior to coarsely guide the iterative merging of superpoints (Yin et al. 2024; Nguyen et al. 2024). In such rough methods, the interplay between semantic and geometric information is not deeply exploited and leads to over-segmentation of object instances. To solve this problem, our unique insight is that leveraging reliable 3D semantics can enhance 3D segmentation and improve the robustness of instance aggregation.
Figure 6: Qualitative comparison on indoor and outdoor datasets. Our method obtains more accurate segmentation results and significantly reduces over-segmentation or fragmented instances compared to previous method. For additional qualitative results, please refer to the supplementary materials.
@@
Leveraging the occlusion-aware viewing visibility, we lift the filtered 2D masks from each view $t$ onto point cloud. These point sets, aggregated by their shared semantic identity, are termed 3D semantic masks, $\mathcal{M}_{3D,t}=\{M_{t,k}\}_{k=1}^{K_{t}}$. While the 2D semantic consistency derived from vision models has proven effective in prior work, the intrinsic geometric coherence of the 3D point cloud itself is often overlooked. We experimentally find that objects with similar appearances are erroneously grouped into a single instance, despite being spatially distinct in the point cloud (see supplementary, Figure 10). As a solution, we apply HDBSCAN (Campello, Moulavi, and Sander 2013) on each of the 3D semantic masks ($M_{t,k}\in\mathcal{M}_{3D,t}$) which splits them into new clusters based on spatial contiguity. We refer to these dense clusters as semantic-geometric seeds, forming a new collection of fine-grained masks $\mathcal{M}_{seed,t}=\{M^{\prime}_{t,j}\}_{j=1}^{K^{\prime}_{t}}$. This density-based splitting utilizes geometric properties to refine the semantics of the segmentation results, ensuring a fine-grained distinction between instances and mitigating the propagation of ambiguous semantics. The effectiveness of this geometric refinement in resolving semantic ambiguity is significant, as our ablation study shows it single-handedly surpasses existing state-of-the-art methods (see Table 2).
@@
While the preceding density-based splitting excels at purifying seeds by enforcing geometric separation, producing $\mathcal{M}_{seed,t}$, it introduces a new challenge: how to correctly ensemble these over-segmented superpoints into complete object instances. To address this, we introduce a high-dimensional feature guided growing process. The core insight is to leverage the rich semantic context from pre-trained features (Schult et al. 2023), to intelligently bridge the spatial gaps between these semantic-geometric seeds and their corresponding object parts.
@@
where $\text{sim}(\bar{\mathbf{f}}_{t,j},\mathbf{f}_{u})$ is the cosine similarity between the seed’s mean feature and the candidate’s mean feature, and $\text{Overlap}(\cdot,\cdot)$ measures their spatial overlap, calculated as the intersection-over-union (IoU) of their 3D point sets. This multiplicative form ensures that a high affinity is only achieved when both semantic similarity and spatial overlap are strong.
At each step of the expansion, we identify the unassigned neighbor with the highest affinity score and merge it into the seed. This process continues iteratively, always selecting the most promising candidate, until the affinity scores of all remaining neighbors become negligible, indicating that no more relevant parts can be found. By optimizing for the highest affinity, object instances can correctly associate disparate parts of the same object and expand its boundary, ready for the final aggregation stage.
Multi-View Progressive Merging
The single-view growing process generates high-fidelity instance proposals, which are essentially the grown semantic-geometric seeds from $\mathcal{M}_{seed,t}$. However, each proposal is inherently limited to the information available from its originating viewpoint. To reconcile these partial views into complete objects, we employ a progressive, multi-view merging strategy. Inspired by (Yin et al. 2024), we initially merge proposals from different views only if their 3D spatial overlap is very high, ensuring only unambiguous matches are made. Subsequent iterations systematically relax this overlap requirement, allowing fragmented parts of the same large instance to be fused. This hierarchical process robustly assembles the final, complete 3D object instances for the entire scene, effectively consolidating information from multiple views.
Experiments
Experimental Setup
@@
Table 1: Quantitative results compared with conventional clustering methods and diverse 2D-to-3D lifting methods on ScanNet200, ScanNet++ and KITTI-360. ”VFM” denotes the vision foundation models. Ours- and Ours+ represent our method without and with ground-truth (GT) depth, respectively. All 2D-to-3D lifting baselines use GT depth. Best results are in bold (%), and the second-best are underlined.
We evaluate our method on three prevalent 3D scene instance segmentation benchmarks: ScanNet200 (Dai et al. 2017), ScanNet++ (Yeshwanth et al. 2023) for indoor environments, and KITTI-360 (Liao, Xie, and Geiger 2022) for outdoor scenes. Following the evaluation protocols established in (Rozenberszki, Litany, and Dai 2024; Schult et al. 2023; Takmaz et al. 2023), we compute AP scores at 50% (AP50) and 25% (AP25) intersection-over-union (IoU) thresholds, as well as the mean AP (mAP) averaged over 50% to 95% IoU thresholds in 5% increments. Implementation details and more experiments are provided in supplementary material.
Analysis Experiments
Quantitative Results
As shown in Table 1, SGS-3D achieves state-of-the-art performance among training-free methods across all benchmarks. The most striking result is on the challenging, depth-less KITTI-360 dataset, where Ours- establishes a remarkable +16.4% mAP lead over the next best competitor, SAI3D. This validates the critical role of our occlusion-aware mapping in robust outdoor scene understanding. On indoor scenes, Ours- also consistently surpasses previous SOTA methods, outperforming Open3DIS on ScanNet200 and SAM2Object on ScanNet++. Furthermore, when provided with GT depth, Ours+ sets a new performance ceiling. It achieves 34.3% mAP on ScanNet200 and 23.7% mAP on ScanNet++, outperforming all training-free baselines and even several training-dependent approaches, which showcases the strong potential of our refinement framework.
Qualitative Results
As shown in Figure 6, our method produces visually superior results with more complete and geometrically precise instances. This enhanced quality stems directly from our framework’s ability to mitigate semantic ambiguity through robust mask filtering and spatial continuity splitting, leading to cleaner boundaries and more coherent object segmentation.
Ablation Study
Table 2 details the contribution of each component. Co-occurrence Mask Filtering (CMF) first boosts mAP to 27.8% by refining initial mask semantics. A key insight comes from then adding Spatial Continuity Splitting (SCP); this variant, relying solely on our purified 3D masks without any superpoint-based growing, achieves an impressive 30.1% mAP. This result alone surpasses state-of-the-art methods like SAM-graph (22.1% mAP), proving the power of our approach in resolving semantic ambiguity at its source. Finally, integrating Feature-Guided Growing (FGG) elevates performance to our full model’s 34.3% mAP, confirming a powerful synergy where our initial purification stages create high-quality seeds for FGG to complete (see Figure 7).
@@
Table 3: Semantic results on ScanNet200. Head, common, and tail categories represent object classes with high, medium, and low frequencies in the dataset, respectively. OpenMask3D (Takmaz et al. 2023) requires supervised training on ScanNet200. In fully zero-shot setting, our method surpass previous SOTA methods by a large margin on all metrics.
@@
Table 4 highlights SGS-3D’s exceptional efficiency and accuracy. Most notably, our method achieves superior accuracy to Open3DIS while using only a quarter of the images (2.5% vs. 10%) and being nearly four times faster. This underscores the remarkable data efficiency and computational performance of our approach. When provided with the same image proportion, SGS-3D’s accuracy advantage widens significantly. This performance gain is also reflected in proposal quality; our method dramatically reduces over-segmentation, cutting the number of predicted instances by nearly 40% compared to Open3DIS (62 vs. 102).
We further evaluated diverse vision foundation models (VFMs, e.g., Cropformer (Qi et al. 2022) and YoloWorld-SAM (Cheng et al. 2024)) to provide initial mask proposals, as shown in Table 6. Our pipeline consistently performs well across different models, demonstrating its robustness. As shown in Table 6, we evaluated the sensitivity of our co-occurrence filtering to the score threshold $c_{m}$. The performance remains stable across a range of values, with the optimal result achieved at 0.2, which we use as our default.
Percentage 0 5 10 30 50 60 70 90
@@
Moreover, we evaluate SGS-3D robustness to occlusion using the patch-drop protocol from (Naseer et al. 2021). Table 7 shows that our method remains highly resilient, maintaining competitive performance even when up to 50% of the key foreground masks are occluded. Performance degrades gracefully under more severe occlusion, confirming the stability of our pipeline.
Figure 8: Open-vocabulary 3D object search. Given a text prompt, our method can accurately locate the corresponding object instances (pink) in a 3D scene, even for rare or uncommon object categories.
@@
The high-quality, class-agnostic instance proposals generated by SGS-3D serve as a powerful foundation for open-set scene understanding. By simply assigning semantic labels to our proposals using an off-the-shelf vision-language model (Zhai et al. 2023), our method can be seamlessly extended to open-vocabulary 3D segmentation (details in supplementary). As shown in Table 3, SGS-3D’s superior performance stems directly from our model’s ability to produce geometrically precise and complete instance masks, which prevents semantic ambiguity and ensures accurate label assignment. This capability enables robust, real-world applications such as interactive, text-based object search within complex 3D environments, as illustrated in Figure 8.

### v2

@@ Abstract
Accurate 3D instance segmentation is crucial for high-quality scene understanding in the 3D vision domain. However, 3D instance segmentation based on 2D-to-3D lifting approaches struggle to produce precise instance-level segmentation, due to accumulated errors introduced during the lifting process from ambiguous semantic guidance and insufficient depth constraints. To tackle these challenges, we propose Splitting and Growing reliable Semantic mask for high-fidelity 3D instance segmentation (SGS-3D), a novel ”split-then-grow” framework that first purifies and splits ambiguous lifted masks using geometric primitives, and then grows them into complete instances within the scene. Unlike existing approaches that directly rely on raw lifted masks and sacrifice segmentation accuracy, SGS-3D serves as a training-free refinement method that jointly fuses semantic and geometric information, enabling effective cooperation between the two levels of representation. Specifically, for semantic guidance, we introduce a mask filtering strategy that leverages the co-occurrence of 3D geometry primitives to identify and remove ambiguous masks, thereby ensuring more reliable semantic consistency with the 3D object instances. For the geometric refinement, we construct fine-grained object instances by exploiting both spatial continuity and high-level features, particularly in the case of semantic ambiguity between distinct objects. Experimental results on ScanNet200, ScanNet++, and KITTI-360 demonstrate that SGS-3D substantially improves segmentation accuracy and robustness against inaccurate masks from pre-trained models, yielding high-fidelity object instances while maintaining strong generalization across diverse indoor and outdoor environments.
@@ Introduction
Arbitrary instances segmentation within 3D scenes constitutes a crucial task in various domains, including autonomous driving, virtual reality, and multi-modality scene understanding. While recent methods (felzenszwalb2004efficient; chen2021hierarchical; vu2022softgroup; liang2021instance; engelmann20203d; lu2023query; schult2023mask3d; yu2023transupr; han2024point; luo2025csfnet) yield impressive segmentation results when trained on specifically annotated datasets (dai2017scannet; armeni20163d; robert2023efficient; yeshwanth2023scannet++; liao2022kitti), their generalization to open-world environments remains limited. In contrast to the inherent difficulties in obtaining extensive 3D annotations, foundation models trained on massive 2D image datasets have shown remarkable performance and strong zero-shot generalization abilities. This observation motivates a promising direction: utilizing 2D pre-trained foundation models for 3D scene perception and interaction.
Currently, two predominant paradigms exist for executing the 3D instance segmentation: feature-based methods and mask-based methods. Feature-based methods (peng2023openscene; takmaz2023openmask3d; ding2023pla; hegde2023clip; yang2024regionplc; lee2024segment; huang2024openins3d) primarily focus on learning robust feature representations from 3D point clouds or by lifting features from 2D modalities (oquab2023dinov2). Instance segmentation is subsequently achieved by classifying, grouping, or decoding these learned features into instance-level masks. However, these methods commonly suffer from inefficient training and error propagation, which arising from ambiguous feature representations stemming from 2D-to-3D lifting or the inherent high-dimensional feature space.
To mitigate these concerns, mask-based methods leverage the semantic masks extracted from input registered images using Segment Anything Model (SAM) (kirillov2023segment) to derive instance-level assignments for each 3D geometric primitive in the 3D scene, consolidating these 3D primitives iteratively based on their projected masks (yang2023sam3d; lu2023ovir; xu2023sampro3d; nguyen2024open3dis; huang2024segment3d). Despite this advancement, the robust propagation of semantic information from these projected masks is often overlooked, with a primary focus on optimizing strategies for merging 3D geometric primitives while neglecting fine-grained geometric cues. This oversight leads to continuous error accumulation during scene segmentation, ultimately yielding imprecise results, particularly for adjacent but separate instances, as il-lustrated in Figure 1. Although some existing methods incorporate depth information from sensors to enhance back-projection accuracy (guo2024sam; yin2024sai3d; zhao2025sam2object), this reliance is problematic for two key reasons: depth sensors inherently struggle on textureless and highly reflective surfaces, and moreover, they are often entirely unavailable in wild scenes, severely compromising segmentation performance.
@@
We first over-segment the point cloud $P$ into a set of $U$ superpoints, $\mathcal{S}=\{s_{u}\}_{u=1}^{U}$, which serve as our fundamental processing primitives. These spatially coherent and geometrically compact groups of points improve computational efficiency and help maintain structural consistency. To adapt to diverse 3D scenarios, we employ method in (dai2017scannet) for indoor mesh inputs and (robert2023efficient) for outdoor unstructured point clouds. Furthermore, for each superpoint $s_{u}$, we assume access to a pre-computed feature vector $\mathbf{f}_{u}\in\mathbb{R}^{D}$, derived from a pre-trained segmentation model (schult2023mask3d), which captures its high-dimensional feature.
@@
In contrast to existing methods that rely on ground-truth depth maps from dedicated sensors, we propose an efficient mapping strategy to compute visibility directly from the point cloud and camera parameters using Z-buffering (ravi2020accelerating). For a given image $I_{t}$ with its associated camera parameters $(\mathbf{K}_{t},\mathbf{T}_{t})$, the entire point cloud $P$ is projected into the camera’s image plane. The transformation $\pi$ maps a homogeneous world point $p_{i}$ to pixel coordinates $(u_{i},v_{i})$ with depth $z_{i}$, according to $z_{i}[u_{i},v_{i},1]^{T}=\mathbf{K}_{t}\mathbf{T}_{t}^{-1}\mathbf{p}_{i}$. To resolve occlusions, a depth buffer $\mathbf{D}_{t}\in\mathbb{R}^{H\times W}$ is initialized with infinite values. The final value at each pixel $(u,v)$ is populated via a highly parallel rasterization process on the GPU, which efficiently solves for the minimum depth over the set of all points projecting to that location:
@@
Mask-based 3D instance segmentation involves generating masks for the target objects based on input images. For each image $I_{t}$ in a scene sequence $I=\{I_{t}\}_{t=1}^{T}$, the objective is to produce a corresponding set of binary masks, $\mathcal{M}_{2D,t}=\{M_{t,k}\}_{k=1}^{K_{t}}$. With the emergence of foundational models like SAM (kirillov2023segment), mask prediction has become highly precise. However, SAM lacks targeted guidance and often segments objects that are not of interest. Therefore, we apply Grounding-DINO (liu2024grounding) to extract box prompts from the images with text prompts. The high-confidence box prompts among them are then used to generate the initial mask predictions.
@@
In fact, the binary masks $\mathcal{M}_{2D,t}$ predicted by the trained 2D vision foundation model exhibit discreteness, which often leads to inaccuracies and inconsistencies in object boundary area predictions across different views. This limitation may increase the instability of the semantic guidance of 2D masks. Unlike Maskclustering (yan2024maskclustering), which purely relies on image pixel consistency between views, we leverage viewing visibility of superpoints to exclusively enhance robustness against inaccurate masks.
@@
Figure 6: Qualitative comparison on indoor and outdoor datasets. Our method obtains more accurate segmentation results and significantly reduces over-segmentation or fragmented instances compared to previous method.
@@
Existing methods often compromise segmentation quality by either directly unprojecting 2D masks onto the point cloud (yang2023sam3d; yan2024maskclustering) or by solely using them as a semantic prior to coarsely guide the iterative merging of superpoints (yin2024sai3d; nguyen2024open3dis). In such rough methods, the interplay between semantic and geometric information is not deeply exploited and leads to over-segmentation of object instances. To solve this problem, our unique insight is that leveraging reliable 3D semantics can enhance 3D segmentation and improve the robustness of instance aggregation.
@@
Leveraging the occlusion-aware viewing visibility, we lift the filtered 2D masks from each view $t$ onto point cloud. These point sets, aggregated by their shared semantic identity, are termed 3D semantic masks, $\mathcal{M}_{3D,t}=\{M_{t,k}\}_{k=1}^{K_{t}}$. While the 2D semantic consistency derived from vision models has proven effective in prior work, the intrinsic geometric coherence of the 3D point cloud itself is often overlooked. We experimentally find that objects with similar appearances are erroneously grouped into a single instance, despite being spatially distinct in the point cloud (see supplementary, Figure 10). As a solution, we apply HDBSCAN (campello2013density) on each of the 3D semantic masks ($M_{t,k}\in\mathcal{M}_{3D,t}$) which splits them into new clusters based on spatial contiguity. We refer to these dense clusters as semantic-geometric seeds, forming a new collection of fine-grained masks $\mathcal{M}_{seed,t}=\{M^{\prime}_{t,j}\}_{j=1}^{K^{\prime}_{t}}$. This density-based splitting utilizes geometric properties to refine the semantics of the segmentation results, ensuring a fine-grained distinction between instances and mitigating the propagation of ambiguous semantics. The effectiveness of this geometric refinement in resolving semantic ambiguity is significant, as our ablation study shows it single-handedly surpasses existing state-of-the-art methods (see Table 2).
@@
While the preceding density-based splitting excels at purifying seeds, producing $\mathcal{M}_{seed,t}$, the challenge is ensembling these over-segmented superpoints into complete instances. To address this, we introduce a feature-guided growing process. We leverage rich pre-trained features (schult2023mask3d) to bridge the spatial gaps between these semantic-geometric seeds and their corresponding object parts.
@@
where $\text{sim}(\bar{\mathbf{f}}_{t,j},\mathbf{f}_{u})$ is the cosine similarity between the seed’s mean feature and the candidate’s mean feature, and $\text{Overlap}(\cdot,\cdot)$ measures their spatial overlap, calculated as the IoU of their 3D point sets. This multiplicative form ensures that a high affinity is only achieved when both semantic similarity and spatial overlap are strong.
@@
Table 1: Quantitative results compared with conventional clustering methods and diverse 2D-to-3D lifting methods on ScanNet200, ScanNet++ and KITTI-360. ”VFM” denotes the vision foundation models. Ours- and Ours+ represent our method without and with ground-truth (GT) depth, respectively. N/A indicates results are not applicable as KITTI-360 lacks the required GT depth. All 2D-to-3D lifting baselines use GT depth in indoor datasets. Best results are in bold (%), and the second-best are underlined.
@@
At each expansion step, we merge the unassigned neighbor with the highest affinity score into the seed. This iterative process continues until the affinity scores of all remaining neighbors become negligible. By optimizing for the highest affinity, object instances can correctly associate disparate parts and expand their boundaries for the final aggregation stage.
Multi-View Progressive Merging
The single-view growing process generates high-fidelity instance proposals, which are essentially the grown semantic-geometric seeds from $\mathcal{M}_{seed,t}$. However, each proposal is inherently limited to the information available from its originating viewpoint. To reconcile these partial views into complete objects, we employ a progressive, multi-view merging strategy. Inspired by (yin2024sai3d), we initially merge proposals from different views only if their 3D spatial overlap is very high, ensuring only unambiguous matches are made. Subsequent iterations systematically relax this overlap requirement, allowing fragmented parts of the same large instance to be fused. This hierarchical process robustly assembles the final, complete 3D object instances for the entire scene, effectively consolidating information from multiple views.
Experiments
Experimental Setup
We evaluate our method on three prevalent benchmarks: ScanNet200 (dai2017scannet), ScanNet++ (yeshwanth2023scannet++) for indoor scenes, and KITTI-360 (liao2022kitti) for outdoor environments. Following established protocols (rozenberszki2024unscene3d; schult2023mask3d; takmaz2023openmask3d), we compute AP scores at 50% (AP50) and 25% (AP25) intersection-over-union (IoU) thresholds, as well as the mean AP (mAP) averaged over 50% to 95% IoU thresholds in 5% increments.
@@
Analysis Experiments
Quantitative Results
As shown in Table 1, SGS-3D achieves SOTA performance among training-free methods. On the challenging, depth-less KITTI-360 dataset, Ours- establishes a remarkable +16.4% mAP lead over the next best competitor, SAI3D, validating our method’s robustness in outdoor scenes. On indoor benchmarks, Ours- also consistently surpasses prior SOTA methods on ScanNet200 and ScanNet++. Furthermore, with GT depth, Ours+ sets a new performance ceiling (e.g., 34.3% mAP on ScanNet200), outperforming all training-free and even several training-dependent approaches, showcasing our framework’s strong potential.
@@
Table 3: Semantic results on ScanNet200. Head, common, and tail categories represent object classes with high, medium, and low frequencies in the dataset, respectively. OpenMask3D (takmaz2023openmask3d) and OpenIns3D (huang2024openins3d) requires supervised training on ScanNet200. In fully supervised and zero-shot setting, our method surpass previous SOTA methods.
@@
Qualitative Results
Figure 6 shows our method produces visually superior results with more complete and geometrically precise instances. This stems from our framework’s ability to mitigate semantic ambiguity through mask filtering and spatial splitting, yielding cleaner and more coherent segmentations.
Ablation Studies
Table 2 details the contribution of each component. CMF provides an initial boost. Critically, adding SCP without growing achieves 30.1% mAP, already surpassing SOTA methods like SAM-graph and proving the power of our semantic purification. Finally, integrating FGG elevates performance to our full model’s 34.3% mAP, confirming a powerful synergy where our purification stages create high-quality seeds for FGG to complete (Figure 7).
@@
Table 4 highlights SGS-3D’s exceptional efficiency. Our method achieves superior accuracy to Open3DIS while using only 2.5% of the images (vs. 10%) and being nearly four times faster. It also significantly reduces over-segmentation by  40%.
Our pipeline demonstrates robustness across diverse vision foundation models (VFMs, Table 6) (qi2022high; cheng2024yolo) and is insensitive to the co-occurrence threshold $c_{m}$ within a reasonable range (Table 6). As shown in Table 7, using a protocol inspired by (naseer2021intriguing), our method is also resilient to occlusion, maintaining strong performance even when up to 50% of foreground masks are removed.
Percentage (%) 0 5 10 30 50 60 70 90
@@
Our high-quality, class-agnostic proposals form a strong basis for open-set understanding. By labeling them with the off-the-shelf vision-language models (radford2021learning; zhai2023sigmoid), we extend our method to open-vocabulary 3D segmentation. As shown in Table 3, our geometrically precise instance masks reduce semantic ambiguity, enabling SGS-3D to significantly outperform previous zero-shot methods and support applications like text-based object search within complex 3D environments, as illustrated in Figure 8.
Figure 8: Open-vocabulary 3D object search. Given a text prompt, our method can accurately locate the corresponding object instances (pink) in a 3D scene.